# Non-linear effects on the Cosmological Gravitational Wave Background anisotropies

### v1

@@ 1 Introduction
The non-thermal nature of GWs makes it difficult to obtain the corresponding relation for the anisotropies in the GW energy density. Typically, the graviton distribution function is expanded as a leading isotropic component that is only a function of the comoving momentum plus a first-order anisotropic contribution. Given that General Relativity is a non-linear theory, it could be important to consider beyond first-order contributions to the anisotropy of the CGWB. Assuming the graviton distribution function can be parameterized by a general power-law form, we can connect the initial anisotropy of the CGWB to the one observed today going beyond perturbation theory. The CGWB anisotropies are generated both by the GW production mechanism and by GW propagation through the perturbed universe. The initial anisotropy depends on the specific mechanisms that sources GWs. We focus on the GWB generated by the quantum vacuum fluctuations of the metric during inflation. In this case, it is possible to obtain a non-perturbative generalization of initial conditions for the CGWB on large scales by recasting the large-scale scalar and small-scale tensor perturbations in such a way as to obtain a new metric, which describes, on large scales, a homogeneous and isotropic Universe. Such a rescaling was employed for the CMB [25, 26] to account for second-order anisotropies and to compute the non-linear generalizations of the Sachs-Wolfe (SW) effect [27]. A similar formalism was extended to include also the Integrated Sachs-Wolfe (ISW) effect in [28].
The non-linearity of the CGWB anisotropies would also affect the higher-order correlation functions. Using the path-integral approach, we compute the 3-point correlation function (and corresponding bispectrum) of the GW energy density, which in the case of GWB generated by quantum tensor fluctuations during inflation with a scale-invariant power spectrum and negligible primordial non-Gaussianity will be exact and hold at any order in perturbation theory.
@@ 2 The non-perturbative Sachs-Wolfe effect for the CGWB
The total GW energy density perturbation includes the intrinsic inhomogeneity of the GW production mechanism and the contribution from the propagation of the gravitons through the large-scale perturbations of the metric, which is provided by the ratio of the comoving momenta at the departure and arrival points.
@@ 2.1 Initial conditions
where $\Phi(\eta,\vec{x})$ and $\Psi(\eta,\vec{x})$ represent the scalar perturbations and $\gamma_{ij}$ are the transverse-traceless tensor fluctuations. Beyond linear order scalar, vector and tensor perturbations are coupled and vector modes could be present. However, we can neglect vector perturbations, since linear vector perturbations are not generated in standard scenarios of inflation as well as in other early universe mechanisms that generate cosmological perturbations, while at higher orders they will contain gradients of scalar modes and thus they can be neglected on large scales.
In the shortwave approximation, it is possible to split tensor perturbations into the small-scale modes that we identify as GWs $h_{ij}$ and the large-scale tensor perturbations $H_{ij}$ [30, 31]. The wavelength of GWs is much smaller than the scales over which the background perturbations vary and hence $\gamma_{ij}=h_{ij}+H_{ij}$. We neglect the large-scale tensor perturbations at all orders, since already at linear order the primordial tensor spectrum is suppressed with respect to the scalar one [32].
@@
$\begin{split}\tilde{a}&=e^{-\Phi_{\ell}}\\ d\tilde{\eta}&=e^{\Psi_{\ell}+\Phi_{\ell}}d\eta\\ \tilde{h}_{ij}&=e^{2\Phi_{\ell}}h_{ij}\,.\end{split}$ (2.14)
The small-scale tensor modes also have to be redefined to avoid mixing of the scalar and tensor perturbations in the definition of GWs. From now on we drop the subscript $\ell$ for simplicity and refer only to superhorizon modes. The new metric then describes a homogeneous and isotropic Universe on large scales plus a modulated small-scale (i.e high-frequency) tensor mode  [^1] Equivalently we can say that the new metric contains also a large-scale tensor mode that arise from a large-scale scalar mode modulating a small-scale (high-frequency) tensor mode, that in fact cannot be reabsorbed into a FLRW metric.
$ds^{2}=\tilde{a}^{2}(\tilde{\eta})\left[d\tilde{\eta}^{2}+(\delta_{ij}+\tilde{h}_{ij})dx^{i}dx^{j}\right]\,.$ (2.15)
@@
$\delta_{\rm GW,\,in}=\frac{e^{-2\Psi+4\Phi}}{\langle e^{-2\Psi+4\Phi}\rangle}-1=\frac{e^{-\Gamma_{\rm in}(n_{\rm gwb}-4)}}{\langle e^{-\Gamma_{\rm in}(n_{\rm gwb}-4)}\rangle}-1\,.$ (2.21)
@@ 3 The CGWB bispectrum
The non-perturbative SW effect can be expressed in terms of the comoving curvature perturbation $\zeta$. This will allow us to relate the anisotropy of the CGWB today directly to the initial conditions determined at the end of inflation. Assuming that all fluids have a definite equation of state, from the continuity equation we derive the following quantity 222 We assume that adiabaticity holds only for standard radiation and matter perturbations and the GW contribution to the total energy density is neglibigle. [35, 36]
@@
Since we are interested in computing the 3-point correlation function of the GW energy density perturbation, we can write
@@
$\begin{split}W^{(3)}(x_{1},x_{2},x_{3})&=\langle(A\,e^{-\zeta_{1}}-1)(A\,e^{-\zeta_{2}}-1)(A\,e^{-\zeta}-1)\rangle_{\rm conn}=\\ &e^{\langle\zeta_{1}\zeta_{2}\rangle+\langle\zeta_{2}\zeta_{3}\rangle+\langle\zeta_{1}\zeta_{3}\rangle}-e^{\langle\zeta_{1}\zeta_{2}\rangle}-e^{\langle\zeta_{2}\zeta_{3}\rangle}-e^{\langle\zeta_{1}\zeta_{3}\rangle}+2\,.\end{split}$ (3.12)
@@ 4 Conclusion
An important aspect of the non-linear effects is the deviation of the primordial perturbations from Gaussian statistics. On large scales, the initial conditions and the SW effect provides the leading contribution to the CGWB anisotropy, however the late ISW effect is also significant. Conversely, effects such as gravitational lensing and Shapiro time-delay become important on smaller scales. The non-linear CGWB anisotropies can affect higher-order correlation functions. The bispectrum of the CGWB is a powerful tool to measure non-Gaussianity on cosmological scales. Besides the primordial contribution generated during or right after inflation, the non-Gaussianity of the CGWB is caused by the non-linear evolution of gravitational potentials. Using the expression for the CGWB anisotropy in terms of the comoving curvature perturbation, we have computed non-perturbatively the 3-point correlation function of the GW energy density. An important result of this work is that in the case of CGWB generated by quantum tensor fluctuations during inflation with the scale-invariant power spectrum and negligible primordial non-Gaussianity, the kernel $\mathcal{K}$, which contains further non-Gaussian contributions, disappears from the expression of the GW energy density perturbation. This will make the SW effect of the CGWB energy density completely insensitive of the possible presence of gravitational slip (i.e., $\Phi\neq\Psi$) even beyond the squeezed limit (as opposed to the adiabatic initial conditions) as arising from e.g. the anisotropic stress produced by collisionless relativistic particles like neutrinos or by modifications of gravity. The effects coming from the presence of anisotropic stress will instead remain in the ISW terms. Finally we have accounted also for (local) primordial non-Gaussianity, providing the perturbative expression for the effective non-linearity parameter $f_{\rm NL}$ characterizing the quadratic non-linearity in the CGWB anisotropies.

### v2

@@ 1 Introduction
The non-thermal nature of GWs makes it difficult to obtain the corresponding relation for the anisotropies in the GW energy density. Typically, the graviton distribution function is expanded as a leading isotropic component that is only a function of the comoving momentum plus a first-order anisotropic contribution. Given that General Relativity is a non-linear theory, it could be important to consider beyond first-order contributions to the anisotropy of the CGWB. Assuming the graviton distribution function can be parameterized by a general power-law form, we can connect the initial anisotropy of the CGWB to the one observed today going beyond perturbation theory. The CGWB anisotropies are generated both by the GW production mechanism and by the GW propagation through the perturbed universe. The initial anisotropy depends on the specific mechanisms that source GWs. We focus on the GWB generated by the quantum vacuum fluctuations of the metric during inflation. In this case, it is possible to obtain a non-perturbative generalization of initial conditions for the CGWB on large scales by recasting the large-scale scalar and small-scale tensor perturbations in such a way as to obtain a new metric, which describes, on large scales, a homogeneous and isotropic universe. Such a rescaling was employed for the CMB [25, 26] to account for second-order anisotropies and to compute the non-linear generalizations of the Sachs-Wolfe (SW) effect [27]. A similar formalism was also extended to include the integrated Sachs-Wolfe effect (ISW) in [28]. We consider a flat spacetime, however a non-zero spatial curvature could enhance the non-linear effects on the CGWB anisotropies [29].
The non-linearity of the CGWB anisotropies would also affect higher-order correlation functions. Using the path-integral approach, we compute the 3-point correlation function (and corresponding bispectrum) of the GW energy density, which in the case of GWB generated by quantum tensor fluctuations during inflation with a scale-invariant power spectrum and negligible primordial non-Gaussianity will be exact and hold at any order in perturbation theory.
@@ 2 The non-perturbative Sachs-Wolfe effect for the CGWB
The total GW energy density perturbation includes the intrinsic inhomogeneity of the GW production mechanism and the contribution from the propagation of gravitons through the large-scale perturbations of the metric, which is provided by the ratio of the comoving momenta at the departure and arrival points.
@@ 2.1 Initial conditions
where $\Phi(\eta,\vec{x})$ and $\Psi(\eta,\vec{x})$ represent the scalar perturbations and $\gamma_{ij}$ are the transverse-traceless tensor fluctuations. Beyond linear order, scalar, vector and tensor perturbations are coupled,, and vector modes could be present. However, we can neglect vector perturbations, since linear vector perturbations are not generated in standard scenarios of inflation as well as in other early universe mechanisms that generate cosmological perturbations, while at higher orders they will contain gradients of scalar modes and thus they can be neglected on large scales.
In the shortwave approximation, it is possible to split tensor perturbations into the small-scale modes that we identify as GWs $h_{ij}$ and the large-scale tensor perturbations $H_{ij}$ [31, 32]. The wavelength of GWs is much smaller than the scales on which the background perturbations vary, and hence $\gamma_{ij}=h_{ij}+H_{ij}$. We neglect the large-scale tensor perturbations at all orders, since already at linear order the primordial tensor spectrum is suppressed with respect to the scalar one [33].
@@
$\begin{split}\tilde{a}&=e^{-\Phi_{\ell}}a\\ d\tilde{\eta}&=e^{\Psi_{\ell}+\Phi_{\ell}}d\eta\\ \tilde{h}_{ij}&=e^{2\Phi_{\ell}}h_{ij}\,.\end{split}$ (2.14)
The small-scale tensor modes also have to be redefined to avoid mixing of the scalar and tensor perturbations in the definition of GWs. From now on, we drop the subscript $\ell$ for simplicity and refer only to superhorizon modes. The new metric then describes a homogeneous and isotropic universe on large scales plus a modulated small-scale (i.e high-frequency) tensor mode  [^1] Equivalently we can say that the new metric contains also a large-scale tensor mode that arise from a large-scale scalar mode modulating a small-scale (high-frequency) tensor mode, that in fact cannot be reabsorbed into a FLRW metric.
$ds^{2}=\tilde{a}^{2}(\tilde{\eta})\left[-d\tilde{\eta}^{2}+(\delta_{ij}+\tilde{h}_{ij})dx^{i}dx^{j}\right]\,.$ (2.15)
@@
$\delta_{\rm GW,\,\rm in}=\frac{e^{-2\Psi+4\Phi}}{\langle e^{-2\Psi+4\Phi}\rangle}-1=\frac{e^{-\Gamma_{\rm in}(n_{\rm gwb}-4)}}{\langle e^{-\Gamma_{\rm in}(n_{\rm gwb}-4)}\rangle}-1\,.$ (2.21)
@@ 3 The CGWB bispectrum
The non-perturbative initial conditions and SW effect can be expressed in terms of the comoving curvature perturbation $\zeta$. This will allow us to relate the anisotropy of the CGWB today directly to the initial conditions determined at the end of inflation. Assuming that all fluids have a definite equation of state, from the continuity equation we derive the following quantity 222 We assume that adiabaticity holds only for standard radiation and matter perturbations and the GW contribution to the total energy density is neglibigle. [36, 37]
@@
While in the CMB case, the non-local contribution to large-scale anisotropy disappears only in the squeezed limit, for the inflationary CGWB instead this is valid for every configuration. The full non-linearity is contained in the curvature perturbation. Since we are interested in computing the 3-point correlation function of the GW energy density perturbation, we can write
@@
$\begin{split}W^{(3)}(x_{1},x_{2},x_{3})&=\langle(A\,e^{-\zeta_{1}}-1)(A\,e^{-\zeta_{2}}-1)(A\,e^{-\zeta_{3}}-1)\rangle_{\rm conn}=\\ &e^{\langle\zeta_{1}\zeta_{2}\rangle+\langle\zeta_{2}\zeta_{3}\rangle+\langle\zeta_{1}\zeta_{3}\rangle}-e^{\langle\zeta_{1}\zeta_{2}\rangle}-e^{\langle\zeta_{2}\zeta_{3}\rangle}-e^{\langle\zeta_{1}\zeta_{3}\rangle}+2\,.\end{split}$ (3.12)
@@ 4 Conclusion
An important aspect of the non-linear effects is the deviation of the primordial perturbations from Gaussian statistics. On large scales, the initial conditions and the SW effect provides the leading contribution to the CGWB anisotropy, however the late ISW effect is also significant. Conversely, effects such as gravitational lensing and Shapiro time-delay become important on smaller scales. The non-linear CGWB anisotropies can affect higher-order correlation functions. The bispectrum of the CGWB is a powerful tool to measure non-Gaussianity on cosmological scales. Besides the primordial contribution generated during or right after inflation, the non-Gaussianity of the CGWB is caused by the non-linear evolution of gravitational potentials. Using the expression for the CGWB anisotropy in terms of the comoving curvature perturbation, we have computed non-perturbatively the 3-point correlation function of the GW energy density. An important result of this work is that in the case of CGWB generated by quantum tensor fluctuations during inflation with the scale-invariant power spectrum and negligible primordial non-Gaussianity, the contribution from the kernel $\mathcal{K}$, which contains further non-Gaussian contributions, vanishes from the expression of the GW energy density perturbation. This will make the SW effect of the CGWB energy density completely insensitive of the possible presence of gravitational slip (i.e., $\Phi\neq\Psi$) even beyond the squeezed limit (as opposed to the adiabatic initial conditions) as arising from e.g. the anisotropic stress produced by collisionless relativistic particles like neutrinos or by modifications of gravity. The effects coming from the presence of anisotropic stress will instead remain in the ISW terms. Finally, we have accounted also for (local) primordial non-Gaussianity, providing the perturbative expression for the effective non-linearity parameter $f_{\rm NL}$ characterizing the quadratic non-linearity in the CGWB anisotropies.